find the slope and length of each side of the following triangles to determine the classification
a= (-3,1) b=(2,3) c=(4,3)
1
Expert's answer
2012-11-20T08:36:19-0500
|ab|=sqrt(25+4)=sqrt(29)
|ac|=sqrt(49+4)=sqrt(53)
|bc|=sqrt(4+0)=2
slope:
ab: arctan(2/5)
ac: arctan(2/7)
bc: arctan(0/2)=0